#25b2 please help, will rate, thanks

The index of refraction for violet and blue light in fused quartz are respectively 1.468 and 1.462. The index of refraction for violet and blue light in polystyrene are respectively 1.506and 1.499. Based on this information, answer the following.

(a) What is the percent difference between the speed of violet light and blue light in fused quartz?
%

(b) What is this percent difference in polystyrene?
%

(c) Which material is more dispersive?

fused quartzpolystyrene
Answer #1

a) Case I: Fused Quartz

V_violet = c/1.468 = 0.681c and V_blue = c/1.462 = 0.684c

where c is the speed of light in vacuum.

Hence, percent difference between the speed = (V_blue - V_violet)*100/c ~ 0.28%

b) Case 2: Polystyrene

V_violet = c/1.506 = 0.664c and V_blue = c/1.499 = 0.667c

where c is the speed of light in vacuum.

Hence, percent difference between the speed = (V_blue - V_violet)*100/c ~ 0.31%

c) Hence, polystyrene is more dispersive than fused quartz because as we saw above that % difference in the speed is more in polystyrene than fused quartz, so the dispersion of white light occurs more in polystyrene and hence more dispersive.